OpenStudy (anonymous):

Solve for x. 2x +3 / 12 = 2x/8 .

pooja195 (pooja195):

Start by amplifying both sides :)

pooja195 (pooja195):

3 / 12 2/8

pooja195 (pooja195):

What can both be simplified to?

pooja195 (pooja195):

1/4 ...........

OpenStudy (anonymous):

oops ment that

pooja195 (pooja195):

lol its ok :) anyways next step \[\huge 2x +1/4 = 1/4x\] subtract 1.4 from both sides \[\huge 2x +1/4 -1/4= 1/4x-1/4\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no its not

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ugh

OpenStudy (kyanthedoodle):

*shoves life cereal down your throat*

OpenStudy (anonymous):

'' gasps''

pooja195 (pooja195):

@whitsbfabulous can you solve what i had above?

pooja195 (pooja195):

\[\huge~2x+1/4−1/4=1/4x−1/4\]

OpenStudy (anonymous):

.......

OpenStudy (anonymous):

i dont know its confusing

pooja195 (pooja195):

Maybe @Nnesha can help :0

OpenStudy (anonymous):

ok

Nnesha (nnesha):

\[\frac{ 2x+3 }{ 12 } = \frac{ 2x }{ 8 }\] or \[2x + \frac{ 3 }{ 12}=\frac{ 2x }{ 8 }\] which one is it ?

OpenStudy (anonymous):

first one

Nnesha (nnesha):

alright cross multiply

Nnesha (nnesha):

 like this